OpenStudy (anonymous):

What is -2 3/8 expression as an improper fraction?

OpenStudy (anonymous):

thanks

OpenStudy (pawanyadav):

-2 3/8=-(2×8+3)/8=-19/8

OpenStudy (anonymous):

-19/8

OpenStudy (pawanyadav):

You are confused

OpenStudy (anonymous):

no, i get it i was typing too fast i'm not confused